# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

### v1

@@ 4.4.1 $a$ appears three times  – $\mathscr{P}(2o,1)$
Without of loss of generality, assume that $a$ appears in $\{a,d,e\}$ in addition to appearing in $\{a,b\}$ and $\{a,c\}$ where neither $d$ nor $e$ are in $\{b,c\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 n1n-1 t1t-1 s2s-2 $\frac{3}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 1 $\frac{1}{4}$
@@
4.4.2 Either $b$ or $c$ appears four times  – $(\neg\bigwedge_{j=1}^{1}\mathscr{P}(2o,j))\land\mathscr{P}(20,2)$
Without of loss of generality, assume that $b$ appears in $\{a,b\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{b,h,i\}$ where none of $\{d,e\}$, $\{f,g\}$ and $\{h,i\}$ is equal to $\{a,c\}$.
@@ 4.4.2.1 Case 1:
There is no element common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 4.4.2.2 Case 2:
There is an element common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$. Let $d=f=h$
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
@@
4.4.3 One of $b$ or $c$ appears only once  – $(\neg\bigvee_{j=1}^{2}\mathscr{P}(2o,j))\land\mathscr{P}(2o,3)$
Assume that $b$ appears only once.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 $\frac{7}{12}$
0 11 $c=1$ n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
@@
4.4.4 There are two sets containing both $b$ and $c$  – $(\neg\bigvee_{j=1}^{3}\mathscr{P}(2o,j))\land\mathscr{P}(2o,4)$
Without of loss of generality, assume we have two sets $\{b,c,d\}$ and $\{b,c,e\}$ where $d\neq e$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 $d=e=1$ n5n-5 t3t-3 s4s-4 $\frac{1}{8}$
11 0 11 n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
11 11 0 n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 $\frac{7}{24}$
@@
4.4.5 There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$  – $(\neg\bigvee_{j=1}^{4}\mathscr{P}(2o,j))\land\mathscr{P}(2o,5)$
Without of loss of generality, assume we have a set $\{b,c,d\}$. Furthermore, assume that there is another set $\{b,e,f\}$ where neither $e$ nor $f$ is equal to $c$ and a set $\{c,g,h\}$ where neither $g$ nor $h$ is equal to $b$. We consider two cases:
4.4.5.1 $d\not\in\{e,f,g,h\}$:
$b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 4.4.5.1 $d\not\in\{e,f,g,h\}$:
4.4.5.2 $d\in\{e,f,g,h\}$:
Assume that $d=e$. In this case, we have the sets $\{a,b\},\{a,c\},\{b,c,d\}$, $\{b,d,f\}$ and $\{c,g,h\}$.
$b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 0 $a=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{5}{12}$
4.4.6 There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$   – $(\neg\bigvee_{j=1}^{5}\mathscr{P}(2o,j))\land\mathscr{P}(2o,6)$
Assume that there are sets $\{b,d,e\}$, $\{c,d,e\}$, and $\{b,f,g\}$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 $2o$ $\frac{5}{12}$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
4.4.7 For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$  – $(\neg\bigvee_{j=1}^{6}\mathscr{P}(2o,j))\land\mathscr{P}(2o,7)$
Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$, $\{c,d^{\prime},e^{\prime}\}$ and $\{c,f^{\prime},g^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$ are distinct.
We consider three cases:
4.4.7.1 Case 1:
There is no element common to three or more of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@
4.4.7.2 Case 2:
There is an element common to exactly three of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$. Assume without loss of generality, $d=f=d^{\prime}$ where $d$ is distinct from $f^{\prime}$ and $g^{\prime}$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
@@
4.4.7.3 Case 3:
There is an element common to exactly four of the sets, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$. Assume without loss of generality, $d=f=d^{\prime}=f^{\prime}$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=g^{\prime}=1$ n8n-8 t5t-5 s7s-7 0 $\frac{1}{24}$
11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{6}{24}$
11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{12}$
@@
Total 11
4.4.8 One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element  – $(\neg\bigvee_{j=1}^{7}\mathscr{P}(2o,j))\land\mathscr{P}(2o,8)$
Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{c,d^{\prime},e^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$ are distinct.
We consider two cases:
Case 1: There is no element that is common to the sets, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@
Total $\frac{35}{36}$
Case 2: There is an element that is common to the sets, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$. Assume $d=f=d^{\prime}$
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{6}{24}$
11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 $2o$ $\frac{5}{36}$
11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{35}{36}$
4.4.9 There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and exactly one set that contains both $b$ and $c$  – $(\neg\bigvee_{j=1}^{8}\mathscr{P}(2o,j))\land\mathscr{P}(2o,9)$
Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\}$, and $\{b,c,d\}$ where $d$ is neither $b$ nor $c$. We consider two cases.
4.4.9.1 $d$ appears uniquely:
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 0 $b=c=1,d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
4.4.9.2 $d$ does not appear uniquely:
@@
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total 11
4.4.10 There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and $b$ and $c$ do not appear together  – $(\neg\bigvee_{j=1}^{9}\mathscr{P}(2o,j))\land\mathscr{P}(2o,10)$
Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $\{d,e,f,g\}\cap\{b,c\}=\varnothing$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 4.5 Proof of the Upper bounds for $\Phi_{2d}(s,t)$
We now prove the upper bounds for $3$-set systems which contain two disjoint $2$-sets as stated in Theorem 4.1. Let $\Xi$} be the class of $3$-set system with two disjoint $2$-sets. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 4$ elements with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains two disjoint $2$-sets $\{a,b\}$ and $\{c,d\}$. We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{5}{6})^{2}(\frac{2}{3})^{\frac{s}{2}}3^{t}$ when $s$ is even and by $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$ when $s$ is odd. We cover $\Xi$} based on the following properties.
@@
There is a set containing a pair of elements from $\{a,b\}\times\{c,d\}$.
@@
A pair of elements appears in two sets where the third element is from $\{a,b,c,d\}$.
@@
There exists two sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
@@
If $\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j)$, then $a$ and $b$ must each appear in another set which contains neither $c$ not $d$ and the two sets cannot have more than one common element.
@@
    no pair from $\{a,b\}\times\{c,d\}$ appears in any set.
Since $a$ and $b$ must participate in two sets, let there be two sets $\{a,e,f\}$ and $\{b,g,h\}$ which are distinct from $\{a,b\}$ and $\{c,d\}$. It must be $\{e,f,g,h\}\cap\{a,b,c,d\}=\varnothing$ since no pair from $\{a,b\}\times\{c,d\}$ appears in any set. Moreover, $\{e,f\}\neq\{g,h\}$ since no pair occurs with one of $\{a,b,c,d\}$ more than once. ∎
4.5.1 One of $\{a,b,c,d\}$ appears 3 times  – $\mathscr{P}(2d,1)$
Assume that $\mathbb{F}$ contains an element in $\{a,b,c,d\}$, say $a$, which appears in three sets. Let $\{a,e,f\}$ and $\{a,g,h\}$ be the other two sets in which $a$ appears. Neither $\{e,f\}$ nor $\{g,h\}$ is equal to $\{c,d\}$.
We consider two cases.
4.5.1.1 Case 1:
There is no common element for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$

### v2

@@ 6.4.1 $a$ appears three times  – $\mathscr{P}(2o,1)$
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,1)$. Without loss of generality, assume that $a$ appears in $\{a,d,e\}$ in addition to appearing in $\{a,b\}$ and $\{a,c\}$ where neither $d$ nor $e$ are in $\{b,c\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
11 n1n-1 t1t-1 s2s-2 0 $\frac{3}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
@@
6.4.2 $\mathscr{P}(2o,2)$ – Either $b$ or $c$ appears four times
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,2)$. Without of loss of generality, assume that $b$ appears in $\{a,b\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{b,h,i\}$ where none of $\{d,e\}$, $\{f,g\}$ and $\{h,i\}$ is equal to $\{a,c\}$.
@@ 6.4.2.1 Case 1:
There is no variable common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@ 6.4.2.2 Case 2:
There is a variable common to $\{d,e\}$, $\{f,g\}$, and $\{h,i\}$. Let $d=f=h$
Core variables Forced variables Restricted CNF parameters Inductive bound
$b$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@
6.4.3 $\mathscr{P}(2o,3)$ – One of $b$ or $c$ appears only once
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,3)$. Assume without loss of generality that $b$ appears only once.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 0 $\frac{7}{12}$
0 11 $c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
@@ 6.4.3 $\mathscr{P}(2o,3)$ – One of $b$ or $c$ appears only once
6.4.4 $\mathscr{P}(2o,4)$ – There are two clauses containing both $b$ and $c$
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,4)$. Without of loss of generality, assume we have two clauses $\{b,c,d\}$ and $\{b,c,e\}$ where $d\neq e$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
11 0 0 $d=e=1$ n5n-5 t3t-3 s4s-4 0 $\frac{1}{8}$
11 0 11 n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
11 11 0 n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
@@
6.4.5 $\mathscr{P}(2o,5)$ – There is exactly one clause in which both $b$ and $c$ appear together and exactly two other clauses, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,5)$. Without of loss of generality, assume we have a clause $\{b,c,d\}$. Furthermore, assume that there is another clause $\{b,e,f\}$ where neither $e$ nor $f$ is equal to $c$ and a clause $\{c,g,h\}$ where neither $g$ nor $h$ is equal to $b$. We consider two cases:
6.4.5.1 Case 1: $d\not\in\{e,f,g,h\}$:
Core variables Forced variables Restricted CNF parameters Inductive bound
$b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@ 6.4.5.1 Case 1: $d\not\in\{e,f,g,h\}$:
6.4.5.2 Case 2: $d\in\{e,f,g,h\}$:
Assume that $d=e$. In this case, we have the clauses $\{a,b\},\{a,c\},\{b,c,d\}$, $\{b,d,f\}$ and $\{c,g,h\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@ 0 $a=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{5}{12}$
6.4.6 $\mathscr{P}(2o,6)$ – There exist clauses $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another clause that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,6)$. Assume that there are clauses $\{b,d,e\}$, $\{c,d,e\}$, and $\{b,f,g\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
11 0 n2n-2 t1t-1 s1s-1 22 $\frac{5}{12}$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
6.4.7 $\mathscr{P}(2o,7)$ – For each $x\in\{b,c\}$, there are exactly two clauses other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,7)$. Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$, $\{c,d^{\prime},e^{\prime}\}$ and $\{c,f^{\prime},g^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$ are distinct. We consider three cases:
6.4.7.1 Case 1: There is no variable common to three or more of the clauses, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@
6.4.7.2 Case 2: There is a variable common to exactly three of the clauses, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
Assume without loss of generality, $d=f=d^{\prime}$ where $d$ is distinct from $f^{\prime}$ and $g^{\prime}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@
6.4.7.3 Case 3: There is a variable common to exactly four of the clauses, $\{d,e\}$, $\{f,g\}$, $\{d^{\prime},e^{\prime}\}$, $\{f^{\prime},g^{\prime}\}$.
Assume without loss of generality, $d=f=d^{\prime}=f^{\prime}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=g^{\prime}=1$ n8n-8 t5t-5 s7s-7 0 $\frac{7}{288}$
11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
11 0 11 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{24}$
11 11 0 n3n-3 t3t-3 s3s-3 $2o$ $\frac{5}{24}$
@@
Total $\frac{283}{288}$
6.4.8 $\mathscr{P}(2o,8)$ – One of $b$ or $c$ appears exactly in three clauses which do not contain the other variable and the other variable appears exactly in two clauses which do not contain the first variable
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,8)$. Assume that $b$ and $c$ only appear in $\{a,b\}$, $\{a,c\}$, $\{b,d,e\}$, $\{b,f,g\}$ and $\{c,d^{\prime},e^{\prime}\}$ where $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$ are distinct. We consider two cases:
6.4.8.1 Case 1: There is no variable that is common to the clauses, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@
11 0 11 $d^{\prime}=e^{\prime}=0$ n5n-5 t2t-2 s1s-1 22 $\frac{5}{36}$
11 11 $c=0$ n3n-3 t2t-2 s3s-3 11 $\frac{1}{4}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{35}{36}$
6.4.8.2 Case 2: There is a variable that is common to the clauses, $\{d,e\}$, $\{f,g\}$, and $\{d^{\prime},e^{\prime}\}$.
Assume $d=f=d^{\prime}$
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
11 0 0 0 $e=g=e^{\prime}=1$ n7n-7 t4t-4 s5s-5 0 $\frac{7}{144}$
11 0 0 11 n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
@@
Total $\frac{47}{48}$
6.4.9 $\mathscr{P}(2o,9)$ – There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and exactly one clause that contains both $b$ and $c$
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,9)$. Assume that $b$ and $c$ only appear in $\{a,b\},\{a,c\}$, and $\{b,c,d\}$ where $d$ is neither $b$ nor $c$. We consider two cases.
6.4.9.1 Case 1: $d$ appears uniquely:
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n$ $t$ $s$ Type Even $s$
@@ 0 $b=c=1,d=0$ n4n-4 t2t-2 s2s-2 0 $\frac{1}{4}$
6.4.9.2 Case 2: $d$ does not appear uniquely:
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n$ $t$ $s$ Type Even $s$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{7}{24}$
Total $\frac{23}{24}$
6.4.10 $\mathscr{P}(2o,10)$ – There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and $b$ and $c$ do not appear together
Assume $\mathbb{F}\in{\Xi}_{2o}(s,t)$ satisfies $\mathscr{P}(2o,10)$. Assume without loss of generality that $b$ and $c$ only appear in $\{a,b\},\{a,c\},\{b,d,e\}$ and $\{c,f,g\}$ where $\{d,e,f,g\}\cap\{b,c\}=\varnothing$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@ 6.5 Proof of the Upper bounds for $\Phi_{2d}(s,t)$
We now prove the upper bounds for monotone $3$-CNFs which contain two disjoint $2$-clauses as stated in Theorem 4.1. Let ${\Xi}_{2d}(s,t)$ be the class of monotone $3$-CNFs over $n\geq 4$ variables with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains two disjoint $2$-clauses $\{a,b\}$ and $\{c,d\}$. We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{5}{6})^{2}(\frac{2}{3})^{\frac{s}{2}}3^{t}$ when $s$ is even and by $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$ when $s$ is odd. We cover ${\Xi}_{2d}(s,t)$ with the following properties.
@@
There is a clause containing a pair of variables from $\{a,b\}\times\{c,d\}$.
@@
A pair of variables appears in two clauses where the third variable is from $\{a,b,c,d\}$.
@@
There exists two clauses $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
@@
If $\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j)$, then $a$ and $b$ must each appear in another clause which contains neither $c$ nor $d$ and the two clauses cannot have more than one common variable.
@@
    no pair from $\{a,b\}\times\{c,d\}$ appears in any clause.
Since $a$ and $b$ must participate in two clauses, let there be two clauses $\{a,e,f\}$ and $\{b,g,h\}$ which are distinct from $\{a,b\}$ and $\{c,d\}$. It must be $\{e,f,g,h\}\cap\{a,b,c,d\}=\varnothing$ since no pair from $\{a,b\}\times\{c,d\}$ appears in any clause. Moreover, $\{e,f\}\neq\{g,h\}$ since no pair occurs with one of $\{a,b,c,d\}$ more than once. ∎
6.5.1 $\mathscr{P}(2d,1)$ – One of $\{a,b,c,d\}$ appears 3 times
Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,1)$. Assume that $\mathbb{F}$ contains an variable in $\{a,b,c,d\}$, say $a$, which appears in three clauses. Let $\{a,e,f\}$ and $\{a,g,h\}$ be the other two clauses in which $a$ appears. Neither $\{e,f\}$ nor $\{g,h\}$ is equal to $\{c,d\}$. We consider two cases
6.5.1.1 Case 1: There is no common variable for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$